X=acos(theta) y=basin(theta) show that dy(square) /dx(square) =-b(power4)/a(square)*y(cube)

Grade:12th pass

1 Answers

Nishant Vora IIT Patna
7 years ago
x =a cosϴ and y = b sinϴ.................(1)

hence x’ = -asinϴ and y’ = b cosϴ

x” = -a cosϴ
y” = -bsinϴ

d^2y/dx^2 = y”/x” = a/b tan ϴ

from eqn 1, tanϴ = bx /ay

hence put the values now